OpenStudy (anonymous):

integral from pi to pi sin^4x cos^5x dx how do it evaluate

OpenStudy (mathmate):

from -pi to pi?

OpenStudy (anonymous):

no both postive

OpenStudy (mathmate):

We know the integral from pi to pi is zero. (same limits) Any integral with lower limit equal to upper limit is zero.

OpenStudy (anonymous):

ok so since sin^4x cos^5x dx. this would conculd to be tan^2x?

OpenStudy (mathmate):

The indefinite integral can be found by substitution: u=sin(x) du = cos(x)dx and use sin^2(x)+cos^2(x)=1 to convert sin^4(x)cos^4(x) =sin^4(x)(1-sin^2(x))^2 =u^4(1-u^2)^2 So I=integral(u^4(1-u^2)^2du And you'll soon find that the integral with limits -pi to pi, -pi to 0, 0 to pi all give 0.

OpenStudy (mathmate):

To make it more understandable, \[\int\limits \sin^4(x)\cos^5(x)dx \ = \ \int\limits u^4(1-u^2)^2du\] with\[u=\sin(x), \ du=\cos(x)dx\]